The prior probabilities for events A1 and A2 are P(A1) = 0.30 and P(A2) = 0.40. It is also known that P(A1 ∩ A2) = 0. Suppose P(B | A1) = 0.20 and P(B | A2) = 0.05. If needed, round your answers to three decimal digits. (a) Are A1 and A2 mutually exclusive? Explain your answer. The input in the box below will not be graded, but may be reviewed and considered by your instructor. (b) Compute P(A1 ∩ B) and P(A2 ∩ B). P(A1 ∩ B) = P(A2 ∩ B) = (c) Compute P(B). P(B) = (d) Apply Bayes’ theorem to compute P(A1 | B) and P(A2 | B). P(A1 | B) = P(A2 | B) =
Answer #1

a)

Yes as P(A1 ∩ A2) = 0

b)

P(A1 n B)=P(A1)*P(B|A1)=03*0.2=0.06

P(A2 n B)=P(A2)*P(B|A2)=0.4*0.05=0.02

c)

P(B)=P(A1 n B)+P(A2 n B)=0.06+0.02=0.08

d) P(A1 | B) =P(A1 n B)/P(B)=0.06/0.08=0.75

P(A2 | B) =0.02/0.08=0.25